# neurips_2024.bib

### v4

@@ Abstract
In this work, we propose a novel adversarial defence mechanism for image classification – Carso – blending the paradigms of adversarial training and adversarial purification in a synergistic robustness-enhancing way. The method builds upon an adversarially-trained classifier, and learns to map its internal representation associated with a potentially perturbed input onto a distribution of tentative clean reconstructions. Multiple samples from such distribution are classified by the same adversarially-trained model, and an aggregation of its outputs finally constitutes the robust prediction of interest. Experimental evaluation by a well-established benchmark of strong adaptive attacks, across different image datasets, shows that Carso is able to defend itself against adaptive end-to-end white-box attacks devised for stochastic defences. Paying a modest clean accuracy toll, our method improves by a significant margin the state-of-the-art for CIFAR-10, CIFAR-100, and TinyImageNet-200 $\ell_{\infty}$ robust classification accuracy against AutoAttack.
@@ 1 Introduction
Vulnerability to adversarial attacks [Biggio2013Evasion, Szegedy2014Intriguing]i.e. the presence of inputs, usually crafted on purpose, capable of catastrophically altering the behaviour of high-dimensional models [Bortolussi2018Intrinsic] – constitutes a major hurdle towards ensuring the compliance of deep learning systems with the behaviour expected by modellers and users, and their adoption in safety-critical scenarios or tightly-regulated environments. This is particularly true for adversarially-perturbed inputs, where a norm-constrained perturbation – often hardly detectable by human inspection [Qin2019Imperceptible, Ballet2019Imperceptible] – is added to an otherwise legitimate input, with the intention of eliciting an anomalous response [Kurakin2018Adversarial].
@@
Given the widespread nature of the issue [Ilyas2019Adversarial], and the serious concerns raised about the safety and reliability of data-learnt models in the lack of an appropriate mitigation [Biggio2017Wild], adversarial attacks have been extensively studied. Yet, obtaining generally robust machine learning (ML) systems remains a longstanding issue, and a major open challenge.
@@
Research in the field has been driven by two opposing, yet complementary, efforts. On the one hand, the study of failure modes in existing models and defences, with the goal of understanding their origin and developing stronger attacks with varying degrees of knowledge and control over the target system [Szegedy2014Intriguing, Goodfellow2015Explaining, MoosaviDezfooli2016DeepFool, Tramer2020OnAdaptive]. On the other hand, the construction of increasingly capable defence mechanisms. Although alternatives have been explored [Cisse2017Parseval, Tramer2018ensemble, Carbone2020Robustness, Zhang2021MEMO], most of the latter is based on adequately leveraging adversarial training [Goodfellow2015Explaining, Madry2018Towards, Tramer2019Multiple, Rebuffi2021Data, Gowal2021Improving, Jia2022LASAT, Singh2023Revisiting, Wang2023Better, Cui2023Decoupled, Peng2023Robust], i.e. training a ML model on a dataset composed of (or enriched with) adversarially-perturbed inputs associated with their correct, pre-perturbation labels. In fact, adversarial training has been the only technique capable of consistently providing an acceptable level of defence [Gowal2020Uncovering], while still incrementally improving up to the current state-of-the-art [Cui2023Decoupled, Peng2023Robust].
@@
Another defensive approach is that of adversarial purification [Shi2021Online, Yoon2021AdvPur], where a generative model is used – similarly to denoising – to recover a perturbation-free version of the input before classification is performed. Nonetheless, such attempts have generally fallen short of expectations due to inherent limitations of the generative models used in early attempts [Nie2022DiffPure], or due to decreases in robust accuracy [^1] The test set accuracy of the frozen-weights trained classifier – computed on a dataset entirely composed of adversarially-perturbed examples generated against that specific model. when attacked end-to-end [Gu2015Towards] – resulting in subpar robustness if the defensive structure is known to the adversary [Tramer2020OnAdaptive]. More recently, the rise of diffusion-based generative models [Huang2021Variational] and their use for purification have enabled more successful results of this kind [Nie2022DiffPure, Chen2023Robust] – although at the cost of much longer training and inference times, and a much brittler robustness evaluation [Chen2023Robust, Lee2024Robust].
@@
In this work, we design a novel adversarial defence for supervised image classification, dubbed Carso (i.e., Counter-Adversarial Recall of Synthetic Observations). The approach relies on an adversarially-trained classifier (called hereinafter simply the classifier), endowed with a stochastic generative model (called hereinafter the purifier). Upon classification of a potentially-perturbed input, the latter learns to generate – from the tensor [^2] Which we call internal representation. of (pre)activations registered at neuron level in the former – samples from a distribution of plausible, perturbation-free reconstructions. At inference time, some of these samples are classified by the very same classifier, and the original input is robustly labelled by aggregating its many outputs. This method – to the best of our knowledge the first attempt to organically merge the adversarial training and purification paradigms – avoids the vulnerability pitfalls typical of the mere stacking of a purifier and a classifier [Gu2015Towards], while still being able to take advantage of independent incremental improvements to adversarial training or generative modelling.
@@
An empirical assessment [^3] Implementation of the method and code for the experiments (based on PyTorch [Paszke2019PyTorch], AdverTorch [Ding2019AdverTorch], and ebtorch [Ballarin2024Ebtorch]) can be found at: https://github.com/emaballarin/CARSO. of the defence in the $\ell_{\infty}$ white-box setting is provided, using a conditional [Sohn2015Learning, Yan2016Attribute2Image] variational autoencoder [Kingma2014Auto, Rezende2014Stochastic] as the purifier and existing state-of-the-art adversarially pre-trained models as classifiers. Such choices are meant to give existing approaches – and the adversary attacking our architecture end-to-end as part of the assessment – the strongest advantage possible. Yet, in all scenarios considered, Carso improves significantly the robustness of the pre-trained classifier – even against attacks specifically devised to fool stochastic defences like ours. Remarkably, with a modest clean accuracy toll, our method improves by a significant margin the current state-of-the-art for Cifar-10 [Krizhevsky2009LearningML], Cifar-100 [Krizhevsky2009LearningML], and TinyImageNet-200 [Chrabaszcz2017Downsampled] $\ell_{\infty}$ robust classification accuracy against AutoAttack [Croce2020AutoAttack].
@@
  • The description of Carso, a novel adversarial defence method synergistically blending adversarial training and adversarial purification;
@@
  • Experimental assessment of the method, against standardised benchmark adversarial attacks – showing higher robust accuracy w.r.t. to existing state-of-the-art adversarial training and purification approaches.
The rest of the manuscript is structured as follows. In section 2 we provide an overview of selected contributions in the fields of adversarial training and purification-based defences – with focus on image classification. In section 3, a deeper analysis is given of two integral parts of our experimental assessment: Pgd adversarial training and conditional variational autoencoders. Section 4 is devoted to the intuition behind Carso, its architectural description, and the relevant technical details that allow it to work. Section 5 contains details about the experimental setup, results, comments, and limitations. Section 6 concludes the paper and outlines directions of future development.
@@ 2 Related work
The idea of training a model on adversarially-generated examples as a way to make it more robust can be traced back to the very beginning of research in the area. The seminal [Szegedy2014Intriguing] proposes to perform training on a mixed collection of clean and adversarial data, generated beforehand.
The introduction of Fgsm [Goodfellow2015Explaining] enables the efficient generation of adversarial examples along the training, with a single normalised gradient step. Its iterative counterpart Pgd [Madry2018Towards] – discussed in section 3 and Appendix A – significantly improves the effectiveness of adversarial examples produced, making it still the de facto standard for the synthesis of adversarial training inputs [Gowal2020Uncovering]. Further incremental improvements have also been developed, some focused specifically on robustness assessment (e.g. adaptive-stepsize variants, as in [Croce2020AutoAttack]).
The most recent adversarial training protocols further rely on synthetic data to increase the numerosity of training datapoints [Gowal2021Improving, Rebuffi2021Data, Wang2023Better, Cui2023Decoupled, Peng2023Robust], and adopt adjusted loss functions to balance robustness and accuracy [Zhang2019Theoretically] or generally foster the learning process [Cui2023Decoupled]. The entire model architecture may also be tuned specifically for the sake of robustness enhancement [Peng2023Robust]. At least some of such ingredients are often required to reach the current state-of-the-art in robust accuracy via adversarial training.
@@
Amongst the first attempts of purification-based adversarial defence, [Gu2015Towards] investigates the use of denoising autoencoders [Vincent2008Extracting] to recover examples free from adversarial perturbations. Despite its effectiveness in the denoising task, the method may indeed increase the vulnerability of the system when attacks are generated against it end-to-end. The contextually proposed improvement adds a smoothness penalty to the reconstruction loss, partially mitigating such downside [Gu2015Towards]. Similar in spirit, [Liao2018Defense] tackles the issue by computing the reconstruction loss between the last-layers representations of the frozen-weights attacked classifier, respectively receiving, as input, the clean and the tentatively denoised example.
In [Samangouei2018DefenseGAN], Generative Adversarial Networks (GANs) [Goodfellow2014Generative] learnt on clean data are used at inference time to find a plausible synthetic example – close to the perturbed input – belonging to the unperturbed data manifold. Despite encouraging results, the delicate training process of GANs and the existence of known failure modes [Zhang2018Convergence] limit the applicability of the method. More recently, a similar approach [Hill2021Stochastic] employing energy-based models [LeCun2006Tutorial] suffered from poor sample quality [Nie2022DiffPure].
Purification approaches based on (conditional) variational autoencoders include [Hwang2019PuVAE] and [Shi2021Online]. Very recently, a technique combining variational manifold learning with a test-time iterative purification procedure has also been proposed [Yang2024Adversarial].
Finally, already-mentioned techniques relying on score- [Yoon2021AdvPur] and diffusion- based [Nie2022DiffPure, Chen2023Robust] models have also been developed, with generally favourable results – often balanced in practice by longer training and inference times, and a much more fragile robustness assessment [Chen2023Robust, Lee2024Robust].
@@ 3 Preliminaries
The task of finding model parameters robust to adversarial perturbations is framed by [Madry2018Towards] as a min-max optimisation problem seeking to minimise adversarial risk. The inner optimisation (i.e., the generation of worst-case adversarial examples) is solved by an iterative algorithm – Projected Gradient Descent – interleaving gradient ascent steps in input space with the eventual projection on the shell of an $\epsilon$-ball centred around an input datapoint, thus imposing a perturbation strength constraint.
@@
Variational autoencoders (VAEs) [Kingma2014Auto, Rezende2014Stochastic] allow the learning from data of approximate generative latent-variable models of the form $p(\bm{x},\bm{z})=p(\bm{x}\operatorname*{|}\bm{z})p(\bm{z})$, whose likelihood and posterior are approximately parameterised by deep artificial neural networks (ANNs). The problem is cast as the maximisation of a variational lower bound.
@@
Conditional Variational Autoencoders [Sohn2015Learning, Yan2016Attribute2Image] extend VAEs by attaching a conditioning tensor $\bm{c}$ – expressing specific characteristics of each example – to both $\bm{x}$ and $\bm{z}$ during training. This allows the learning of a decoder model capable of conditional data generation.
@@ 4.1 Architectural overview and principle of operation
There are no specific requirements for the classifier, whose training is completely independent of the use of the model as part of Carso. However, training it adversarially improves significantly the clean accuracy of the overall system, allowing it to benefit from established adversarial training techniques.
@@
In fact, the decoder of a conditional VAE allows for exact algorithmic differentiability [Baydin2018Automatic] w.r.t. its conditioning set, thus averting the need for backward-pass approximation [Athalye2018Obfuscated] in generating end-to-end adversarial attacks against the entire system, and preventing (un)intentional robustness by gradient obfuscation [Athalye2018Obfuscated]. The same cannot be said [Chen2023Robust] for more capable and modern purification models, such as those based e.g. on diffusive processes, whose robustness assessment is still in the process of being understood [Lee2024Robust].
@@
Upon completion of the training process, the encoder network may be discarded as it will not be used for inference.
@@
Remarkably, the only link between the initial potentially-perturbed input and the resulting purified reconstructions (and thus the predicted class) is through the internal representation of the classifier, which serves as a featurisation of the original input. The whole process is exactly differentiable end-to-end, and the only potential hurdle to the generation of adversarial attacks against the entire system is the stochastic nature of the decoding – which is easily tackled by Expectation over Transformation [Athalye2018Synthesizing].
@@ 4.2 A first-principles justification
If we consider a trained ANN classifier, subject to a successful adversarial attack by means of a slightly perturbed example, we observe that – both in terms of $\ell_{p}$ magnitude and human perception – a small variation on the input side of the network is amplified to a significant amount on the output side, thanks to the layerwise processing by the model. Given the deterministic nature of such processing at inference time, we speculate that the trace obtained by sequentially collecting the (pre)activation values within the network, along the forward pass, constitutes a richer characterisation of such an amplification process compared to the knowledge of the input alone. Indeed, as we do, it is possible to learn a direct mapping from such featurisation of the input, to a distribution of possible perturbation-free input reconstructions – taking advantage of such characterisation.

### v5

# and Aggregation Improves Adversarial Robustness

@@ Abstract
In this work, we propose a novel adversarial defence mechanism for image classification – Carso – blending the paradigms of adversarial training and adversarial purification in a synergistic robustness-enhancing way. The method builds upon an adversarially-trained classifier, and learns to map its internal representation associated with a potentially perturbed input onto a distribution of tentative clean reconstructions. Multiple samples from such distribution are classified by the same adversarially-trained model, and a carefully chosen aggregation of its outputs finally constitutes the robust prediction of interest. Experimental evaluation by a well-established benchmark of strong adaptive attacks, across different image datasets, shows that Carso is able to defend itself against adaptive end-to-end white-box attacks devised for stochastic defences. Paying a modest clean accuracy toll, our method improves by a significant margin the state-of-the-art for Cifar-10, Cifar-100, and TinyImageNet-200 $\ell_{\infty}$ robust classification accuracy against AutoAttack.
@@ 1 Introduction
Vulnerability to adversarial attacks (Biggio et al., 2013; Szegedy et al., 2014)i.e. the presence of inputs, usually crafted on purpose, capable of catastrophically altering the behaviour of high-dimensional models (Bortolussi & Sanguinetti, 2018) – constitutes a major hurdle towards ensuring the compliance of deep learning systems with the behaviour expected by modellers and users, and their adoption in safety-critical scenarios or tightly-regulated environments. This is particularly true for adversarially-perturbed inputs, where a norm-constrained perturbation – often hardly detectable by human inspection (Qin et al., 2019; Ballet et al., 2019) – is added to an otherwise legitimate input, with the intention of eliciting an anomalous response (Kurakin et al., 2018).
@@
Given the widespread nature of the issue (Ilyas et al., 2019), and the serious concerns raised about the safety and reliability of data-learnt models in the lack of an appropriate mitigation (Biggio & Roli, 2018), adversarial attacks have been extensively studied. However, obtaining generally robust machine learning (ML) systems remains a longstanding issue, and a major open challenge.
@@
Research in the field has been driven by two opposing, yet complementary, efforts. On the one hand, the study of failure modes in existing models and defences, with the goal of understanding their origin and developing stronger attacks with varying degrees of knowledge and control over the target system (Szegedy et al., 2014; Goodfellow et al., 2015; Moosavi-Dezfooli et al., 2016; Tramèr et al., 2020). On the other hand, the construction of increasingly capable defence mechanisms. Although alternatives have been explored (Cisse et al., 2017; Tramèr et al., 2018; Carbone et al., 2020; Zhang et al., 2022), most of the latter is based on adequately leveraging adversarial training (Goodfellow et al., 2015; Madry et al., 2018; Tramèr & Boneh, 2019; Rebuffi et al., 2021; Gowal et al., 2021; Jia et al., 2022; Singh et al., 2023; Wang et al., 2023; Cui et al., 2023; Peng et al., 2023), i.e. training a ML model on a dataset composed of (or enriched with) adversarially-perturbed inputs associated with their correct, pre-perturbation labels. In fact, adversarial training has been the only technique capable of consistently providing an acceptable level of defence (Gowal et al., 2020), while still incrementally improving up to the current state-of-the-art (Cui et al., 2023; Peng et al., 2023; Bartoldson et al., 2024).
@@
Another defensive approach is that of adversarial purification (Shi et al., 2021; Yoon et al., 2021), where a generative model is used – similarly to denoising – to recover a perturbation-free version of the input before classification is performed. Nonetheless, such attempts have generally fallen short of expectations due to inherent limitations of the generative models used in early attempts (Nie et al., 2022), or due to decreases in robust accuracy [^1] The test set accuracy of the frozen-weights trained classifier – computed on a dataset entirely composed of adversarially-perturbed examples generated against that specific model. when attacked end-to-end (Gu & Rigazio, 2015) – resulting in subpar robustness if the defensive structure is known to the adversary (Tramèr et al., 2020). More recently, the rise of diffusion-based generative models (Huang et al., 2021) and their use for purification have enabled more successful results of this kind (Nie et al., 2022; Chen et al., 2023) – although at the cost of longer training and inference times, and a much brittler robustness evaluation (Chen et al., 2023; Lee & Kim, 2024).
@@
In this work, we design a novel adversarial defence for supervised image classification, dubbed Carso (i.e., Counter-Adversarial Recall of Synthetic Observations). The approach relies on an adversarially-trained classifier (called hereinafter simply the classifier), endowed with a stochastic generative model (called hereinafter the purifier). Upon classification of a potentially-perturbed input, the latter learns to generate – from the tensor [^2] Which we call internal representation. of (pre)activations registered at neuron level in the former – samples from a distribution of plausible, perturbation-free reconstructions. At inference time, some of these samples are classified by the very same classifier, and the original input is robustly labelled by aggregating its many outputs in the form of a normalised doubly-exponential logit product. This method – to the best of our knowledge the first attempt to organically merge the adversarial training and purification paradigms – avoids the vulnerability pitfalls typical of the mere stacking of a purifier and a classifier (Gu & Rigazio, 2015), while still being able to take advantage of independent incremental improvements to adversarial training or generative modelling.
@@
An empirical assessment [^3] Implementation of the method and code for the experiments (based on PyTorch (Paszke et al., 2019), AdverTorch (Ding et al., 2019), and ebtorch (Ballarin, 2024)) can be found at: https://github.com/emaballarin/CARSO. of the defence in the $\ell_{\infty}$ white-box setting is provided, using a conditional (Sohn et al., 2015; Yan et al., 2016) variational autoencoder (Kingma & Welling, 2014; Rezende et al., 2014) as the purifier and existing state-of-the-art adversarially pre-trained models as classifiers. Such choices are meant to give existing approaches – and the adversary attacking our architecture end-to-end as part of the assessment – the strongest advantage possible. Yet, in all scenarios considered, Carso improves significantly the robustness of the pre-trained classifier – even against attacks specifically devised to fool stochastic defences like ours. Remarkably, with a modest clean accuracy toll, our method improves by a significant margin the current state-of-the-art for Cifar-10 (Krizhevsky, 2009), Cifar-100 (Krizhevsky, 2009), and TinyImageNet-200 (Chrabaszcz et al., 2017) $\ell_{\infty}$ robust classification accuracy against AutoAttack (Croce & Hein, 2020a).
@@
  • The description of Carso, a novel adversarial defence method synergistically blending adversarial training and adversarial purification, thanks to representation-conditional purification and a dedicated robust aggregation strategy;
@@
  • An experimental assessment of the method, against standardised benchmark adversarial attacks – showing higher robust accuracy w.r.t. to existing state-of-the-art adversarial training and purification approaches.
The rest of the manuscript is structured as follows. In section 2 we provide an overview of selected contributions in the fields of adversarial training and purification-based defences – with focus on image classification. In section 3, a deeper analysis is given of two integral parts of our experimental assessment: Pgd adversarial training and conditional variational autoencoders. Section 4 is devoted to the intuition behind Carso, its architectural description, and the relevant technical details that allow it to work effectively. Section 5 contains details about the experimental setup, results, comments, and limitations. Section 6 concludes the paper and outlines directions of future development.
@@ 2 Related work
The idea of training a model on adversarially-generated examples as a way to make it more robust can be traced back to the very beginning of research in the area. In their seminal work, Szegedy et al. (2014) propose to perform training on a mixed collection of clean and adversarial data, generated beforehand.
The introduction of Fgsm (Goodfellow et al., 2015) enables the efficient generation of adversarial examples along the training, with a single normalised gradient step. Its iterative counterpart Pgd (Madry et al., 2018) – discussed in section 3 and Appendix A – significantly improves the effectiveness of the adversarial examples produced, making it still the de facto standard for the synthesis of adversarial training inputs (Gowal et al., 2020). Further incremental improvements have also been developed, some focused specifically on robustness assessment (e.g. stepsize-adaptive variants, as by Croce & Hein (2020a)).
The most recent adversarial training protocols further rely on synthetic data to increase the numerosity of training datapoints (Gowal et al., 2021; Rebuffi et al., 2021; Wang et al., 2023; Cui et al., 2023; Peng et al., 2023; Bartoldson et al., 2024), and adopt adjusted loss functions to balance robustness and accuracy (Zhang et al., 2019a) or generally foster the learning process (Cui et al., 2023). The entire model architecture may also be tuned specifically for the sake of robustness enhancement (Peng et al., 2023). At least some of such ingredients are often required to reach the current state-of-the-art in robust accuracy via adversarial training.
@@
Amongst the first attempts of purification-based adversarial defence, Gu & Rigazio (2015) investigate the use of denoising autoencoders (Vincent et al., 2008) to recover examples free from adversarial perturbations. Despite its effectiveness in the denoising task, the method may indeed increase the vulnerability of the system when attacks are generated against it end-to-end. The contextually proposed improvement adds a smoothness penalty to the reconstruction loss, partially mitigating such downside (Gu & Rigazio, 2015). Similar in spirit, Liao et al. (2018) tackle the issue by computing the reconstruction loss between the last-layers representations of the frozen-weights attacked classifier, respectively receiving, as input, the clean and the tentatively denoised example.
In the work by Samangouei et al. (2018), Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) learnt on clean data are used at inference time to find a plausible synthetic example – close to the perturbed input – belonging to the unperturbed data manifold. Despite encouraging results, the delicate training process of GANs and the existence of known failure modes (Zhang et al., 2018) limit the applicability of the method. More recently, a similar approach (Hill et al., 2021) employing energy-based models (LeCun et al., 2006) suffered from poor sample quality (Nie et al., 2022).
Purification approaches based on (conditional) variational autoencoders include the works by Hwang et al. (2019) and Shi et al. (2021). Very recently, a technique combining variational manifold learning with a test-time iterative purification procedure has also been proposed (Yang et al., 2024).
Finally, already-mentioned techniques relying on score- (Yoon et al., 2021) and diffusion- based (Nie et al., 2022; Chen et al., 2023) models have also been developed, with generally favourable results – often balanced in practice by longer training and inference times, and a much more fragile robustness assessment (Chen et al., 2023; Lee & Kim, 2024).
@@ 3 Preliminaries
The task of finding model parameters robust to adversarial perturbations is framed by Madry et al. (2018) as a min-max optimisation problem seeking to minimise adversarial risk. The inner optimisation (i.e., the generation of worst-case adversarial examples) is solved by an iterative algorithm – Projected Gradient Descent – interleaving gradient ascent steps in input space with the eventual projection on the shell of an $\epsilon$-ball centred around an input datapoint, thus imposing a perturbation strength constraint.
@@
Variational autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) allow the learning from data of approximate generative latent-variable models of the form $p(\bm{x},\bm{z})=p(\bm{x}\operatorname*{|}\bm{z})p(\bm{z})$, whose likelihood and posterior are approximately parameterised by deep artificial neural networks (ANNs). The problem is cast as the maximisation of a variational lower bound.
@@
Conditional Variational Autoencoders (Sohn et al., 2015; Yan et al., 2016) extend VAEs by attaching a conditioning tensor $\bm{c}$ – expressing specific characteristics of each example – to both $\bm{x}$ and $\bm{z}$ during training. This allows the learning of a decoder model capable of conditional data generation.
@@ 4.1 Architectural overview and principle of operation
There are no specific requirements for the classifier, whose training is completely independent of the use of the model as part of Carso. However, training it adversarially significantly improves the robust accuracy of the overall system, also allowing it to benefit from established adversarial training techniques.
@@
In fact, the decoder of a conditional VAE allows for exact algorithmic differentiability (Baydin et al., 2018) w.r.t. its conditioning set, thus averting the need for backward-pass approximation (Athalye et al., 2018a) in generating end-to-end adversarial attacks against the entire system, and preventing (un)intentional robustness by gradient obfuscation (Athalye et al., 2018a). The same cannot be said (Chen et al., 2023) for more capable and modern purification models, such as those based e.g. on diffusive processes, whose robustness assessment is still in the process of being thoroughly understood (Lee & Kim, 2024).
@@
Upon completion of the training process, the encoder network is discarded, as it will not be used for inference.
@@
Remarkably, the only link between the initial potentially-perturbed input and the resulting purified reconstructions (and thus the predicted class) happens through the internal representation of the classifier, which serves as a featurisation of the original input. The whole process is exactly differentiable end-to-end, and the only potential hurdle to the generation of adversarial attacks against the entire system is the stochastic nature of the decoding – which is easily tackled by Expectation over Transformation (Athalye et al., 2018b).
@@ 4.2 A first-principles justification
If we consider a trained ANN classifier, subject to a successful adversarial attack by means of a slightly perturbed example, we observe that – both in terms of $\ell_{p}$ magnitude and human perception (Bartoldson et al., 2024) – a small variation on the input side of the network is amplified to a significant amount on the output side, thanks to the layerwise processing by the model. Given the deterministic nature of such processing at inference time, we speculate that the trace obtained by sequentially collecting the (pre)activation values within the network, along the forward pass, constitutes a richer characterisation of such an amplification process compared to the knowledge of the input alone. Indeed, as we do, it is possible to learn a direct mapping from such featurisation of the input, to a distribution of possible perturbation-free input reconstructions – in a way that takes advantage of such characterisation.